Paulie ordered 250 pens and 250 pencils to sell for a theatre club fundraiser. The pens cost 11 cents more than the pencils. If paulie's total order cost $42.50, find the cost of each pen and pencil.
Answer by fractalier(6550)

Just did this one...
Call the cost of the pencils, x.
Thus the cost of the pens are x + 11 (we can work in cents).
Then we can write
250(x + 11) + 250x = 4250
250x + 2750 + 250x = 4250
500x + 2750 = 4250
500x = 1500
x = 3 cents (pencils)
x+11 = 14 cents (pens)